A car covered a distance of 300 km with a speed of 100 kmph.Find the average speed of the car ?

Grade:6

1 Answers

Vijay
2 years ago
Explanation -
formula for Average Speed = ( total distance ) / ( total time taken ).  ---- (1)
Here Total distance = 300 km ,
total time = ( distance ) / (speed) ;
so we get = 300/100 = 3 hrs.
 
Now Avg. speed = (300) / (3) = 100kmph.
 
required answer is 100kmph.
thanks